OpenStudy (anonymous):

A ball of mass 5 slugs is dropped from a height of 1000ft. Find the limiting velocity of the ball if it encounters a force due to air resistance equal to (-1/2)v.

OpenStudy (dumbcow):

whats limiting velocity? do they just want the velocity function in terms of t

OpenStudy (dumbcow):

Force of gravity is mass* acceleration ma = 5*(-32) = -160 Force of air resistance is -(1/2)v Net force: -160-(-1/2v) = (1/2)v -160 Net acceleration, divide by the mass: a = (1/10)v - 32 Net velocity: **initial velocity is zero** v = (1/10)v*t -32t solve for v: v -(1/10)vt = -32t v(1-(1/10)t) = -32t v = -32t/(1-(1/10)t)

OpenStudy (dumbcow):

If you then take the limit as t goes to infinity. multiply top and bottom by 1/t lim = -32/(1/t -1/10) -> lim 1/t = 0 lim = -32/-1/10 = 32*10 = 320 Then limiting velocity would be 320 ft/sec *If thats what they mean*

OpenStudy (dumbcow):

you may want another opinion, im not as confident in my physics i hope i did this right

OpenStudy (anonymous):

This is correct, I just checked :)

OpenStudy (dumbcow):

ok good to hear :)